OpenStudy (anonymous):

A falcon flew 70 miles in 3 hours. What was the speed of the falcon? speed = distance over time; 1 mile = 1609 meters. A. 10.4 meters per second B. 23.3 meters per second C. 52.2 meters per second D. 93.9 meters per second

OpenStudy (anonymous):

speed= distance/time 1mile = 1609 meters 70 miles ==70*1609 =112630 miles 1 hour = 3600 second 3 hours = 3*3600 =10800 so the speed = 112630/10800 = ???

OpenStudy (anonymous):

@jorge0013